OpenStudy (anonymous):

[2.05] Mike is six less than three times his daughter's age. Mike is 36; how old is his daughter? 10 12 14 16

OpenStudy (anonymous):

14. 36 + 6 = 42. 42 / 3 = 14!

OpenStudy (ybarrap):

@supanick14 Do you understand why ? M = Mike's age. D = Daughter's age. M = \(3\times D - 6 =36 \implies3D = 36+6 \implies D=14 \).

OpenStudy (anonymous):

ya